Calculate the length of the polar curve r=1-sin(theta) on the given interval (0, 2π).


1
Expert's answer
2021-05-05T07:42:20-0400

length of an arc "l" ="\int\limits_a^b \sqrt{{r' (}\theta)^2 + r(\theta)^2 d\theta}"

where "r = 1 -sin\theta"


and "r' =(1-cos )^2"

then "l=" "\int\limits_a^b \sqrt{(1-cos )^2 + (1 -sin\theta)^2}"

where "a =0" and "b= 2\pi"


"l=" 6.283183